At the exit of a toll gate with a single booth, vehicle arrive at random at a rate of 20 vehicles per minute. The service has an average rate of 22 vehicles per minute. Estimate the following:


(a) Average length of queue formed at the toll gate.


(b) Average waiting time of vehicles.


(c) Average time vehicles spent in the system.




1
Expert's answer
2021-10-11T07:01:16-0400

Part a

"Q= \frac{\rho^2}{1- \rho}\\
Q= \frac{(\frac{20}{22})^2}{1- \frac{20}{22}}\\
Q=9.08 vehicles"


Part b

"t= (\frac{1}{t_1}+\frac{1}{t_2})*60\\
t= (\frac{1}{22}+\frac{1}{20})*60\\
t=5.73 sec"


Part c

"t= \frac{2- \rho}{2 \mu (1- \rho)}\\
t= \frac{2- \frac{20}{22}}{2*22 (1-\frac{20}{22})}\\
t=0.27 min"